# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

### v2

@@ 1. INTRODUCTION
Limited necessity for explicit interactions. As observed in Figure 1, most models using only explicit feature interactions achieve AUC performance below 81.3, while DNN achieves an AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNNs, which undoubtedly undermines the necessity of integrating explicit and implicit feature interactions. Therefore, FinalMLP (Mao et al., 2023) attempts to model features implicitly in a dual manner, discarding traditional explicit interaction methods, and thereby achieving state-of-the-art performance.
Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in higher-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two to three network layers, abandoning the explicit capture of effective higher-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times, potentially offsetting the benefits gained from improved model accuracy.
Insufficient and undifferentiated supervision signals. Most models using both explicit and implicit feature interaction methods require a fusion layer to obtain the final prediction (Wang et al., 2017, 2021b; Mao et al., 2023). However, they only use the final prediction to compute the loss, rather than providing appropriate supervision signals for the different methods themselves. This weakens the effectiveness of the supervision signals. Additionally, some studies (Wang et al., 2023b; Liu et al., 2024), such as the CL4CTR in Figure 1, attempt to introduce auxiliary loss to provide extra supervision signals. However, this often introduces additional computational costs and loss balancing hyperparameters, increasing the difficulty of hyperparameter tuning. Therefore, a simple, general, and effective method for computing supervision signals is crucial.
To address the aforementioned limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and the Shallow & Deep Cross Network (SDCNv3), which integrates both low-order and high-order feature interactions while ensuring model interpretability by avoiding the use of DNN for implicit high-order feature interaction modeling. Specifically, we introduce a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and we use a Self-Mask operation to filter noise and reduce the number of parameters in the Cross Network by half. In the fusion layer, we propose a simple yet effective multi-loss balancing strategy and calculation method, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
@@
We introduce a novel feature interaction modeling method, Deep Crossing, which grows exponentially with the number of layers to achieve a genuine deep cross network (DCNv3). This method explicitly captures feature interaction information while using a Self-Mask operation to reduce the number of parameters by half and filter noise.
We propose a model, SDCNv3, that explicitly captures both low-order and high-order feature interactions. Additionally, we introduce a simple and effective multi-loss balancing and calculation method, called Tri-BCE, to ensure that different sub-networks receive appropriate supervision signals.
@@ 2.1. CTR Prediction
Effectively capturing feature interactions has always been one of the key methods for improving CTR prediction models, thus receiving extensive research attention. Traditional methods include LR (Richardson et al., 2007), which captures first-order feature interactions, and FM (Rendle, 2010) and its derivatives (Sun et al., 2021; Xiao et al., 2017; Pan et al., 2018), which capture second-order feature interactions. With the rise of deep learning, several models attempt to use DNN to capture higher-order feature interactions (e.g., PNN (Qu et al., 2016), Wide & Deep (Cheng et al., 2016), DeepFM (Guo et al., 2017), DCNv1 (Wang et al., 2017), DCNv2 (Wang et al., 2021b), and DIN (Zhou et al., 2018)), achieving better performance. Among these, the DCN series models are widely recognized for their effective trade-off between efficiency and performance, gaining significant attention from both academia and industry (Wang et al., 2017, 2021b; Chen et al., 2021; Wang et al., 2023a; Yu et al., 2021; Lian et al., 2018). Most subsequent deep CTR models follow the paradigm established by DCN, integrating explicit and implicit feature interactions.
@@ 2.2. PROBLEM DEFINITION
Further, we can define a CTR sample in the tuple data format: . Variable is an true label for user click behavior:
@@
Implicit feature interaction aims to automatically learn complex non-manually defined data patterns and high-order feature interactions using DNNs. It is characterized by high efficiency, strong performance, and poor interpretability. Additionally, (Rendle et al., 2020) points out the inefficiency of DNNs in learning multiplicative operations.
@@ 3.1. Embedding & Reshape Layer
The input for the CTR prediction task typically consists of , , and , which are multi-field categorical data and are represented using one-hot encoding. Most CTR prediction models utilize an embedding layer to transform them into low-dimensional dense vectors: , where and separately indicate the embedding matrix and the vocabulary size for the -th field, represents the embedding dimension. In our model, to enable the Cross & Masked Vector to share a weight matrix and separately process two different information streams, we reshape the embeddings using the chunk operation to divide them into two different views: . Further, we can get the resulting embedding after reshaping: , where denotes the number of fields, , and denotes first-order features.
@@ 3.2. Shallow & Deep Cross Network v3
where represents the order feature interaction. From Figure 3, we can see that the original Cross Network v2 is essentially a shallow crossing method, achieving linear growth in the order of feature interactions through the stacking of layers. Experiments show that the optimal number of cross layers in DCNv2 is 2 to 3 (Wang et al., 2021b), with most of its performance coming from the DNN responsible for implicit interactions. In contrast, DCNv3 modifies to to achieve exponential growth in the order of feature interactions.
@@
Other mask mechanisms can also be used here, such as random Mask based on the Bernoulli distribution, learnable Mask based on Top-K selection, etc. To ensure our proposed model is simple and effective, we use LayerNorm to perform a straightforward and efficient normal distribution transformation on , ensuring that its output contains approximately 50% zero values to filter out noise, and it reduces the number of parameters by half compared to Cross Network v2.
@@ 3.3. Fusion Layer
Tri-BCE loss calculation and balancing method are shown in Figure 4. We use the widely adopted binary cross-entropy loss (Wang et al., 2023a; Zhu et al., 2023b; Mao et al., 2023) (i.e., Logloss) as both the primary and auxiliary loss for the SDCNv3:
@@
As demonstrated by (Li et al., 2024), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by :
@@ 4.2. Overall Performance (RQ1)
Overall, SDCNv3 achieves the best performance across all six datasets. SDCNv3 shows an average AUC improvement of 0.22% over the strongest baseline model and an average Logloss improvement of 0.11%, both exceeding a statistically significant performance improvement of 0.1%. This demonstrates the effectiveness of SDCNv3. SDCNv3 ranks 1st in CTR prediction on the PapersWithCode benchmarks999https://paperswithcode.com/task/click-through-rate-prediction for the Criteo, KDD12, and KKBox datasets.
@@ 4.3.1. Efficiency Comparison (RQ2)
Explicit CTR models typically use fewer parameters. For instance, LR, FM, FwFM, and AFM have nearly zero non-embedding parameters, while FmFM, CrossNet, CIN, and AutoInt all require fewer than 1M parameters. Notably, parameter count does not always correlate with time complexity. Although CIN uses only 0.57M parameters, its training time per epoch reaches a maximum of 606 seconds, making it unsuitable for practical production environments.
@@
Our proposed DCNv3 and SDCNv3 are the most parameter-efficient models in the DCN series, requiring only 0.78M and 1.37M parameters respectively to achieve SOTA performance. Meanwhile, regarding runtime, DCNv3 consistently outperforms strong baseline models such as FinalMLP, FINAL, DCNv2, and DCNv1. This demonstrates the time-efficient of DCNv3. Although SDCNv3 requires an additional 20 seconds compared to DCNv3 due to the employment of the Tri-BCE loss, it still remains comparable to DCNv2. Notably, the extra computational cost for the loss is only incurred during training and does not affect inference speed in practical applications.
4.3.2. The Interpretability of SDCNv3 (RQ3)
@@ 5. CONCLUSION
This paper introduces the next generation deep cross networks, DCNv3 and SDCNv3. The former explicitly captures feature interaction through an exponentially growing modeling method, and further filters noise signals via the Self-Mask operation, reducing the parameter count by half. The latter builds on DCNv3 by incorporating the shallow cross network, SCNv3, to capture both high-order and low-order feature interactions without relying on the less interpretable DNN. Tri-BCE helps the two sub-networks in SDCNv3 obtain more suitable supervision signals for themselves. Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. Additionally, our proposed models achieve 1st rankings in multiple CTR benchmarks using only explicit feature interactions, breaking the convention that traditional CTR models must integrate implicit feature interactions to improve performance.

### v4

@@ 1. Introduction
Limited necessity for explicit interactions. As observed in Figure 1, most models (e.g., CIN (Lian et al., 2018)) using only explicit feature interactions achieve AUC performance below 81.3, while DNN achieves a better AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNN, which undoubtedly undermines the necessity of integrating explicit and implicit feature interactions. Therefore, FinalMLP (Mao et al., 2023) attempts to model features implicitly in a dual manner, discarding traditional explicit interaction methods, and thereby achieving state-of-the-art performance.
Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in higher-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two or three network layers, abandoning the explicit capture of effective higher-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times, potentially offsetting the benefits gained from improved model accuracy.
Insufficient and undifferentiated supervision signals. Most models using both explicit and implicit feature interaction methods require a fusion layer to obtain the final prediction (Wang et al., 2017, 2021b; Mao et al., 2023). However, they only use the final prediction to compute the loss, rather than providing appropriate supervision signals for the different methods themselves. This weakens the effectiveness of the supervision signals. Additionally, some studies (Wang et al., 2023b; Liu et al., 2024), such as the CL4CTR (Wang et al., 2023b) in Figure 1, attempt to introduce auxiliary loss to provide extra supervision signals. However, this often introduces additional computational costs and loss trade-off hyperparameters, increasing the difficulty of hyperparameter tuning. Therefore, a simple, general, yet effective method for computing supervision signals is crucial.
To address the aforementioned limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and the Shallow & Deep Cross Network (SDCNv3), which integrates both low-order and high-order feature interactions while ensuring model interpretability by avoiding the use of DNN for implicit high-order feature interaction modeling. Specifically, we introduce a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and we design a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we propose a simple yet effective multi-loss trade-off strategy and calculation method, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
@@
We introduce a novel feature interaction modeling method, called Deep Crossing, which grows exponentially with the number of layers to achieve a genuine deep cross network (DCNv3). This method explicitly captures feature interaction information while using a Self-Mask operation to reduce the number of parameters by half and filter noise.
We propose a novel CTR model, called SDCNv3, that explicitly captures both low-order and high-order feature interactions. Additionally, we introduce a simple yet effective multi-loss trade-off and calculation method, called Tri-BCE, to ensure that different sub-networks receive appropriate supervision signals.
@@ 2.1. CTR Prediction
Effectively capturing feature interactions has always been one of the key methods for improving CTR prediction, thus receiving extensive research attention (Tian et al., 2023; Wang et al., 2023a; Zhu et al., 2023b). Traditional methods include LR (Richardson et al., 2007), which captures first-order feature interactions, and FM (Rendle, 2010) and its derivatives (Sun et al., 2021; Xiao et al., 2017; Pan et al., 2018), which capture second-order feature interactions. With the rise of deep learning, several models attempt to use DNN to capture higher-order feature interactions (e.g., PNN (Qu et al., 2016), Wide & Deep (Cheng et al., 2016), DeepFM (Guo et al., 2017), DCNv1 (Wang et al., 2017), DCNv2 (Wang et al., 2021b), and DIN (Zhou et al., 2018)), achieving better performance. Among these, the DCN series models are widely recognized for their effective trade-off between efficiency and performance, gaining significant attention from both academia and industry (Wang et al., 2017, 2021b; Chen et al., 2021; Wang et al., 2023a; Yu et al., 2021; Lian et al., 2018). Most subsequent deep CTR models follow the paradigm established by DCN, integrating explicit and implicit feature interactions.
@@ 2.2. Problem Definition
Further, we can define a CTR input sample in the tuple data format: . Variable is an true label for user click behavior:
@@
Implicit feature interaction aims to automatically learn complex non-manually defined data patterns and high-order feature interactions using DNN. It is characterized by high efficiency, strong performance, and poor interpretability. Additionally, (Rendle et al., 2020) points out the inefficiency of DNN in learning multiplicative operations.
@@ 3.1. Embedding & Reshape Layer
The input of the CTR prediction task, which is multi-field categorical data and is represented using one-hot encoding. Most CTR prediction models (Song et al., 2019; Li et al., 2023; Sang et al., 2024) utilize an embedding layer to transform them into low-dimensional dense vectors: , where and separately indicate the embedding matrix and the vocabulary size for the -th field, represents the embedding dimension. In our model, to enable the Cross & Masked Vector to share a weight matrix and separately process two different information streams, we reshape the embeddings using the chunk operation to divide them into two different views: . Further, we can get the resulting embedding after reshaping: , where denotes the number of fields, , and denotes the first-order features.
@@ 3.2. Shallow & Deep Cross Network v3
where represents the order feature interaction. From Figure 3, we can see that the original Cross Network v2 is essentially a shallow crossing method, achieving linear growth in the order of feature interactions through the stacking of layers. Experiments show that the optimal number of cross layers in DCNv2 is 2 or 3 (Wang et al., 2021b), with most of its performance coming from the DNN responsible for implicit interactions. In contrast, DCNv3 modifies to to achieve exponential growth in the order of feature interactions.
@@
Other mask mechanisms can also be used here, such as random Mask based on the Bernoulli distribution, learnable Mask based on Top-K selection, etc. To ensure our proposed model is simple and effective, we use LayerNorm (Ba et al., 2016) to perform a straightforward and efficient normal distribution transformation on , ensuring that its output contains approximately 50% zero values to filter out noise, and it reduces the number of parameters by half compared to Cross Network v2.
@@ 3.3. Fusion Layer
Tri-BCE loss calculation and trade-off method are shown in Figure 4. We use the widely adopted binary cross-entropy loss (Wang et al., 2023a; Zhu et al., 2023b; Mao et al., 2023) (i.e., Logloss) as both the primary and auxiliary loss for the SDCNv3:
@@
As demonstrated by (Li et al., 2024), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better parameters by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by :
@@ 4.2. Overall Performance (RQ1)
Overall, SDCNv3 achieves the best performance across all six datasets. SDCNv3 shows an average AUC improvement of 0.22% over the strongest baseline model and an average Logloss decrease of 0.11%, both exceeding a statistically significant performance improvement of 0.1%. This demonstrates the effectiveness of SDCNv3. SDCNv3 ranks 1st in CTR prediction on the PapersWithCode benchmarks999https://paperswithcode.com/task/click-through-rate-prediction for the Criteo, KDD12, and KKBox datasets.
@@ 4.3.1. Efficiency Comparison (RQ2)
Explicit CTR models typically use fewer parameters. For instance, LR, FM, FwFM, and AFM have nearly zero non-embedding parameters, while FmFM, CrossNet, CIN, and AutoInt all require fewer than 1M parameters. Notably, parameter count does not always correlate with time complexity. Although CIN uses only 0.57M parameters, its training time per epoch reaches a maximum of 606 seconds, making it unsuitable for practical production environments. FiGNN faces the same issue.
@@
Our proposed DCNv3 and SDCNv3 are the most parameter-efficient models in the DCN series, requiring only 0.78M and 1.37M parameters respectively to achieve SOTA performance. Meanwhile, regarding runtime, DCNv3 consistently outperforms strong baseline models such as FinalMLP, FINAL, DCNv2, and DCNv1. This demonstrates the time-efficient of DCNv3. Although SDCNv3 requires an additional 20 seconds compared to DCNv3 due to the employment of the Tri-BCE loss, it still remains comparable to DCNv2. Notably, the extra computational cost for the loss is only incurred during training and does not affect inference speed in practical applications. This further demonstrates the efficiency of DCNv3 and SDCNv3.
4.3.2. The Interpretability and noise filtering capability of SDCNv3 (RQ3)
@@ 5. Conclusion
This paper introduced the next generation deep cross networks, DCNv3 and SDCNv3. The former explicitly captured feature interaction through an exponentially growing modeling method, called Deep Crossing, and further filtered noise signals via the Self-Mask operation, reducing the parameter count by half. The latter built on DCNv3 by incorporating the shallow cross network, called SCNv3, to capture both high-order and low-order feature interactions without relying on the less interpretable DNN. Tri-BCE helped the two sub-networks in SDCNv3 obtain more suitable supervision signals for themselves. Comprehensive experiments on six datasets demonstrated the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. Additionally, our proposed models achieved 1st rankings in multiple CTR benchmarks using only explicit feature interactions, breaking the convention that traditional CTR models must integrate implicit feature interactions to improve performance.
@@ Acknowledgements.
- Sang et al. (2024) Lei Sang, Honghao Li, Yiwen Zhang, Yi Zhang, and Yun Yang. 2024. AdaGIN: Adaptive Graph Interaction Network for Click-Through Rate Prediction. ACM Transactions on Information Systems (2024).